OpenStudy (anonymous):

Find the average rate of change of the function from x(1) to x(2). f(x) = x^2-3x+6. x-values: x(1) is 1, x(2) is 5.

OpenStudy (anonymous):

average rate of change is integral of function f(x) divided by x(2) - x(1) so integral of f(x) is 2x-3 and from 1 to 5 is (10-3)-(2-3)=8 so 8 / x(2) - x(1) = 8/(5-1)=2...so the answer is 2

OpenStudy (anonymous):

But that's actually wrong.